Please write out the 3 that apply.[47500] Question 6 Select the three correct statements below. Three companies have purchased the pre-diabetes medication. The
Answer #1

Let a= \text{ price of one pill A}~,~b= \text{ Price of one pill B} and C= fixed overhead charges

Then we have to solve the system

{\bf X}{\bf b}={\bf z} Where X and Z are as given. And {\bf b}= \begin{bmatrix} a\\ b \\ c\end{bmatrix}

The augmented matrix for teh above system is

[ {\bf X} : {\bf z}]= \begin{bmatrix} 1000& 800 &1 &:& 45000\\ 800& 1000& 1&:& 46000 \\ 1500 & 500& 1&:& 47500\end{bmatrix} \\ ~~~~~~~~~~~~\sim \begin{bmatrix} 1000& 800 &1 &:& 45000\\ -200& 200& 0&:& 1000 \\ 500 & -300& 0&:& 2500\end{bmatrix}

By applying the transformation R_2=R_2-R_1, R_3=R_3-R_1

Now again apply the transformation R_3=R_3+{3\over 2}R_2 we get

[ {\bf X} : {\bf z}]\sim \begin{bmatrix} 1000& 800 &1 &:& 45000\\ -200& 200& 0&:& 1000 \\ 200 & 0& 0&:& 4000\end{bmatrix} ---------------(1)

From the last row we get

200a=4000 \\ \Rightarrow a=20

So Price of one pill A is $20

From the 2nd row of (1) we get

-200(20)+200b=1000 \\ \Rightarrow b=25

So Price of one pill B is $25

Similarly from row 1st of (1) we get

c=5000

So the fixed overhead charges are $5000

Hence the correct Options are First, Third and Fifth